OpenStudy (anonymous):

solve the system x - 2y + z =4 -x + 4y - 3z =2 2x + y + 10z=10

OpenStudy (anonymous):

What math are you in because you can either put this in a matrix which is faster or you can write it all out by hand

OpenStudy (anonymous):

Do you know matrices?

OpenStudy (anonymous):

nope. i think it has to be the traditional way

OpenStudy (anonymous):

Then you need to solve for each variable one at a time. Do you need help getting started?

OpenStudy (anonymous):

@rshextall yes please! :)

OpenStudy (anonymous):

Then lets start with the first equation and solve for x.

OpenStudy (anonymous):

Just add 2y and subtract z from both sides of the equation \[x-2y+2y+z-z=4+2y-z\] \[x=4+2y-z\]

OpenStudy (anonymous):

Do you see what happened there?

OpenStudy (anonymous):

yup !

OpenStudy (anonymous):

Then you need to plug it into one of the other equations and solve for either y or z

OpenStudy (anonymous):

so plugging it into the 2nd equation mates it 2y-2z=6 ? then?

OpenStudy (anonymous):

Right!!!

OpenStudy (anonymous):

It takes a while but it's not to difficult, but its easy to make simple mistakes like forgetting a minus sign and stuff

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

Then solve for either y or z

OpenStudy (anonymous):

Then you have to both x and the other into the last equation and solve for the last variable

OpenStudy (anonymous):

You can check your answer by plugging your answers for x, y, and z into the first equations. They should work for all 3 equations

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

z=-1 y=2 x=9. yup i got it thanks!

OpenStudy (anonymous):

@rshextall but can you show me how to do the matrix, it's in the next lesson of the book so i think i should learn it too.

OpenStudy (anonymous):

You have to learn the rules of matrices. They're a bit more in depth than we can really get into on here. If you like I can help you with your homework with it when you get it